OpenStudy (anonymous):

use an algebraic equation to find the measures of the two angles described below. Begin by letting x represent the degree measure of the angle's supplement. The measure of the angle is 8 degrees greater than its complement What is the measure of the complement? What is the measure of the other angle? @wolfe8

OpenStudy (ranga):



OpenStudy (anonymous):

i'm so confused here..

OpenStudy (anonymous):

Lol..I thought ranga was gonna answer your question

OpenStudy (ranga):

AOB is the angle in question. The problem says assume x to be the supplement of the angle in question which is DOB. AOB = 180 - x COB is the complement of the angle in question = DOB - DOC = x - 90 The problem says AOB is 8 degrees greater than its complement. AOB = COB + 8 180 - x = x - 90 + 8 Solve for x

OpenStudy (ranga):

Sorry, I had to leave in the middle for a phone call.

OpenStudy (ranga):

180 - x = x - 90 + 8 180 - x = x - 82 180 + 82 = 2x 2x = 262 x = 131 AOB = 180 - 131 = 49 COB = AOB - 8 = 41

OpenStudy (anonymous):

Sorry i was working out the problem

OpenStudy (anonymous):

Thank you for your help. I was stuck on this for a long time

OpenStudy (ranga):

You are welcome.

OpenStudy (ranga):

What is the measure of the complement? 41 degrees What is the measure of the other angle? I am assuming they mean x is the other angle and so it is 131. But you decide what they mean by the other angle. The angle in question is AOB = 49 but they don't ask for it.